OpenStudy (anonymous):

Brothers Jack and Poy do brisk walking from their house to school daily. Jack walks at 9 km/h, and oy walks at 5 km/h. When Jack reaches the school, Poy is 500 m away. How far is the school from the boys' house?

OpenStudy (anonymous):

1.125km

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

Let the time taken by Jack to reach the school be t. the same time will be taken by Poy to travel the distance which is 500 m away from the school. s=vt where s-displacement, v-velocity and t-time. 9t=0.5+5t 4t=o.5 t=0.125 s=9t =9*0.125 =1.125km